OpenStudy (weasel123):

is 2 √x+5+7=4 extraneous?

rebeccaxhawaii (rebeccaxhawaii):

An extraneous solution is a root of a transformed equation that is not a root of the original equation because it was excluded from the domain of the original equation.

OpenStudy (weasel123):

so is it?

OpenStudy (jdoe0001):

\(\bf 2\sqrt{x+5}+7=4?\)

OpenStudy (weasel123):

is 2 √x+5+7=4 extraneous?

OpenStudy (jdoe0001):

can't quite see who's the radicand... so, did you mean \(\bf \bf 2\sqrt{x+5}+7=4?\)